# Dirac Fermions and Topological Phases in Magnetic Topological Insulator Films

### v2

@@ I Introduction
The Hall conductivity carried by a massive or gapless Dirac fermion is discussed generally, with additional symmetry constraints imposed on the Fermi surface for the latter one, for both continuum and lattice models. A direct deduction leads to the result that the Hall conductivities associated with the gapless and gapped Dirac fermions in the TI film are ±$\pm e^{2}/2h$ and 00, respectively, leading to a half quantum mirror Hall effect by 1/21/$2-1/$ 2 - 1 / 2, serving as a metallic partner to the insulating quantum spin Hall effect. A brief proof for the half-quantization of a metallic band structure with considered symmetry constraints over the Fermi surface is also presented. Additionally, a field theoretical deduction for the half quantization, and a discussion on handling the Hall conductivity of a gapless Dirac fermion are provided.
@@
In the weak Zeeman field regime, the topological phases are characterized by focusing on $n=1$ matrix elements affecting the two gapless Dirac fermions near the surface. This framework clarifies the underlying physics behind the Chern insulator, axion insulator, and half QAHE, with symmetric, antisymmetric, or unilateral distribution of Zeeman fields at the surface of the TI film, respectively. The resulting Hall conductance exhibits quantized nature: $1+0$, $0+0$, and $1/2+0$ in units of $e^{2}/h$. Additionally, the mirror layer Chern number in the Chern insulator with symmetrically distributed magnetism is examined, revealing $(1/4)$$(1/2)$$(1/4)$ partition for the non-trivial band and $(c/4)$$(-c/2)$$(c/4)$ with $c\approx 1$ for the trivial band.
@@ II Massless and massive Dirac fermions in a topological insulator film
Figure 1: Schematic momentum dependent mass configurations (upper panel) and corresponding band structure of Dirac fermions (lower panel). The quantization of Hall conductivity is denoted by $N$ for half quantization in a metallic band and $C$ for quantization at the bottom of a gapped band. The colors assigned to the Dirac cones represent the sign of Berry curvature with red for positive and blue for negative. (a) On the left panel, two gapless Dirac fermions are shown, whose masses are zero at low-energy near the Dirac point (assumed to be $k=0$), while non-vanishing at high-energy with opposite signs, which we define as the chirality $\chi$ of a 2D gapless Dirac fermion. Such chirality unambiguously determines the sign of the loop integral of Berry connection around the Fermi surface, consequently determining the sign of half-quantized Hall conductivity. (b) In the middle, two trivially gapped massive Dirac fermions are present, with masses being either positive or negative for all $k$, leading to a sign change of Berry curvature and a totally vanishing Hall conductivity labeled by zero Chern number. (c) On the right panel, two non-trivial gapped Dirac cones are displayed, and the corresponding masses exhibit kink configurations with sign change between low and high-energy areas. Such non-trivial mass configuration indicates overall Berry curvature sign convergence, and leads to a non-vanishing Hall conductivity labelled by an integer Chern number. The non-triviality is also addressed by formally drawing states connecting conduction and valence bands, well-known as chiral edge states for a Chern band under open boundary conditions[13, 14].
@@
for the continuum and lattice model, respectively. Here we adopt a homogeneous in-film-plane parameter set with $a$ and $\lambda_{\parallel}$ as the in plane lattice constant and Fermi velocity, and $\bm{k}=(k_{x},k_{y})$ is the in-film-plane wavevector. Notice that an infinitely direct summed Dirac fermions exist in the continuum model, while there are only $2L_{z}$ species with $L_{z}$ the layer number along opened $z$-direction of the film in the lattice model. For the mainly concerned individual Dirac cone with a single Dirac point at $k=0$, its topological property is revealed based on a general discussion over the nature of its Hall conductivity quantization, as revealed in the schematic diagram Fig. 1. Especially, in the strong TI film with a single Dirac cone at $\Gamma$, aka $k=0$ point, the gapless pair of Dirac fermions carry ±$\pm e^{2}/2h$, as half-quantized Hall conductivity, while the gapped series are all trivial.
@@ II.1 The continuum model
$\varphi^{n}(s)$ $=C\begin{pmatrix}-is\lambda_{\perp}f_{+}^{n}\\ t_{\perp}\eta^{n}f_{-}^{n}\end{pmatrix},\leavevmode\nobreak\ E=m_{n},$ (4a)
$\chi^{n}(s)$ $=C\begin{pmatrix}t_{\perp}\eta^{n}f_{-}^{n}\\ is\lambda_{\perp}f_{+}^{n}\end{pmatrix},\leavevmode\nobreak\ E=-m_{n},$ (4b)
where the dependence on $(\bm{k},z)$ is inherited inside even/odd parity functions $f_{\pm}^{n}(\bm{k},z)$ and real factor $\eta^{n}(\bm{k})$, whose definition can be found in Appendix A. The $k$-dependent eigenvalue of $h(s)$ is represented by ±$\pm m_{n}(\bm{k}),\leavevmode\nobreak\ n=1,2,\cdots$, as a mass term, which can be solved in a closed manner through equations
@@
$\xi_{\alpha}$ $=\sqrt{-\frac{F}{D}+(-1)^{\alpha-1}\frac{\sqrt{R}}{D}},\leavevmode\nobreak\ \alpha=1,2,$ (5b)
@@
Figure 2: The momentum-dependent mass of Dirac fermions in a TI film as a continuum model. The lowest several momentum-dependent $m_{n}(\bm{k})$ along $k_{x}$ solved from closed equations Eq. (5) are presented, while the homogeneous in-plane nature of the model ensures that the asymptotic behavior of $m_{n}(\bm{k})$ is the same as $k\rightarrow\infty$. Here the film thickness $L=(L_{z}+1)c$ with $L_{z}=10$ is chosen here as a TI film with $10$ layers. The index $n$ is assigned such that $|m_{n}|$ increases with $n$. Especially notice the sign-jump behavior that $\mathrm{sgn}(m_{n}(\infty))=(-1)^{n}$. From here on, the model parameters on lattice for numerical calculations and verifications are set as $\lambda_{\parallel}=0.41\leavevmode\nobreak\ $\mathrm{e}\mathrm{V}$$ eV, $\lambda_{\perp}=0.44\leavevmode\nobreak\ $\mathrm{e}\mathrm{V}$$, $t_{\parallel}=0.566\leavevmode\nobreak\ $\mathrm{e}\mathrm{V}$$, $t_{\perp}=0.4\leavevmode\nobreak\ $\mathrm{e}\mathrm{V}$$, $m_{0}=0.28\leavevmode\nobreak\ $\mathrm{e}\mathrm{V}$$, $a=b=1\leavevmode\nobreak\ $\mathrm{n}\mathrm{m}$$, $c=0.5\leavevmode\nobreak\ $\mathrm{n}\mathrm{m}$$ if with no specific indication[54]. Generically, these parameters can be determined through the first-principle calculations, and the specific choice here is for the sake of illustration. This parameter choice makes the bulk 3D TI a strong one with a single Dirac point at $\Gamma$. And for the continuum model discussed here, the substitution $\lambda_{\parallel}\rightarrow\lambda_{\parallel}a$ $\lambda_{\perp}\rightarrow\lambda_{\perp}c$, $t_{\parallel}\rightarrow t_{\parallel}a^{2}$, $t_{\perp}\rightarrow t_{\perp}c^{2}$ should be recognized.
@@
with $\gamma=m_{0}t_{\perp}/\lambda_{\perp}^{2},u=\lambda_{\perp}/2t_{\perp}$. The numerical result is shown in Fig. 3, with excellent agreement between the lowest order approximated gap and that from solving the set of non-linear equations, especially for relatively large $L$. The exponentially decaying tendency is best revealed by the logarithmic absolute value of mass gap at $k=0$, as its center decreases linearly with thickness, while the oscillating nature is revealed by the dips, which will extend to negative infinity at strict gap closing point, and the mass gap will reverse its sign before and after the dip, as shown directly by the $m_{1}(0)-L$ diagram and the inner amplified picture. Since $m_{1}(\infty)=m_{0}(\infty)<0$ is certain, we see that the oscillating behavior of $m_{1}(0)$ with thickness $L$ can drive $m_{1}(\bm{k})$ to share configuration that jumps between the one shown in Fig. 1(b) and (c), i.e., between a trivial band and a band with unit Chern number. Then for an ultra-thin film which owns two copies ±$\pm m_{1}(\bm{k})$ reflected by $\tau_{z}$ in Eq. (7), the $\mathbb{Z}_{2}$ topological index shows jumping behavior between $\mathbb{Z}_{2}=0$ and $\mathbb{Z}_{2}=1$, i.e., between a band insulator and a quantum spin Hall insulator[40, 59, 60, 41, 42]. We will not discuss further about this phenomenon except for giving an explicit $\mathbb{Z}_{2}(L_{z})$ oscillating diagram below in the lattice model subsection shown in Fig. 5. We emphasize here that the exponentially decaying gap will not affect physically observable topological phase, either for an insulating or metallic one, for a TI film with enough thickness.
@@ II.2 The lattice model
where energy and hopping matrices are $\mathcal{M}_{0}=(m_{0}-2\sum_{\mu}t_{\mu})\beta$, $\mathcal{T}_{\mu}=t_{\mu}\beta-i\frac{\lambda_{\mu}}{2}\alpha_{\mu}$, with $l$ and $\mu$ denoting site locations and three spatial directions, while $\{\beta,\alpha_{\mu}\}$ denoting Dirac matrices under standard Dirac representation $\beta=\sigma_{0}\tau_{z},\leavevmode\nobreak\ \alpha_{\mu}=\sigma_{\mu}\tau_{x},$ where Pauli matrices $\sigma_{\mu}$ and $\tau_{\mu}$ represent different degrees of freedom, respectively. For instance, one could choose them to represent spin and pseudo-spin (like orbital) ones. $\Psi_{l}$ represents vectorized Fermionic operator at site $l$. Notice that when adopting a full Fourier transformation upon all three spatial dimensions, i.e., an infinite bulk system, the Hamiltonian is transformed into the standard modified Dirac’s equation[27] on lattice $\mathcal{H}_{TI}=\sum_{\bm{k}}\Psi_{\bm{k}}^{\dagger}H(\bm{k})\Psi_{\bm{k}}$ where
@@
$\sin\xi_{\alpha}=\sqrt{1-\cos\xi_{\alpha}^{2}},\leavevmode\nobreak\ \alpha=1,2.$ (22)
Now, different from the continuum model, the set of equations give $L_{z}$ solutions $m_{n}(\bm{k}),n=1,2,\cdots,L_{z}$ including one surface state and $L_{z}-1$ purely trivial bulk states, if within suitable choice of parameters. This is essentially because now the Dirac equation is put on lattice, and the number of solutions is constrained by finite lattice constants. And the other set of $L_{z}$ masses are just the chiral partners with eigenvalues $-m_{n}(\bm{k})$.
Figure 4: The momentum-dependent mass of Dirac fermions in a TI film on the lattice. Namely, $m_{n}(\bm{k}),\leavevmode\nobreak\ n=1,2,\cdots,L_{z}$ along $k_{x}$ are solved from closed equations Eq. (20) of lattice model with $L_{z}=40$. Again, index $n$ is assigned in the way that $|m_{n}(\pi,\pi)|$ increases with $n$. Especially notice the sign-jump behavior that $\mathrm{sgn}(m_{n}(\pi,\pi))=(-1)^{n}$.
@@
with $\chi=\pm$ labelling the mirror eigenvalue[33]. An example of $m_{n}(k)$ with $L_{z}=40$ is presented in Fig. 4. Among these Dirac fermions, two of them with ±$\pm m_{1}(\bm{k})$ are gapless Dirac cones with their low-energy states localized at top and bottom surfaces, while emerging into the bulk at their high-energy away from Dirac point, and the remaining fermions are all gapped. Notice that the same arguments about the projection as a trivial local unitary transformation and Heaviside Theta function form of the lowest solution (see below) can be made here, as in the continuum model.
@@ III.1.1 Gapless/metallic case
$h_{1,\chi}=\lambda_{\parallel}(\bm{k}\cdot\bm{\sigma})+\chi\Theta(-m_{0}(k))m_{0}(k)\sigma_{z},\leavevmode\nobreak\ \chi=\pm.$ (32)
@@ III.1.2 Insulating case
for a Dirac cone, where $d_{\min}=\min(d(k))$ denotes the bound of the global gap. Clearly, $\sigma_{H}/(e^{2}/h)=0,\pm 1$ appears, notifying trivial or non-trivial phases depending on the relative signs of low and high-energy masses, with the ±$\pm 1$ cases identified as the Chern insulator or equivalently, the quantum anomalous Hall effect. In our equivalent model, one sees from Fig. 2 that all $n\geq 2$ masses contains the same sign, and the corresponding Dirac cones are all trivial. And we come back to the statement that in a TI film, there are two gapless Dirac fermions with opposite half-quantized Hall conductance, while all other bands form paired trivial massive Dirac fermions. The quantized nature of the Hall conductance in insulating system, $\sigma_{H}=-Ce^{2}/h$, is referred to by the famous TKNN theorem[10], with its robustness against continuous non-gap-closing perturbations rooted in the topological nature of $C$ as the Chern invariant[69, 70].

### v3

@@ I Introduction
The Hall conductivity carried by a massive or gapless Dirac fermion is discussed generally, with additional symmetry constraints imposed on the Fermi surface for the latter one, for both continuum and lattice models. A direct deduction leads to the result that the Hall conductivities associated with the gapless and gapped Dirac fermions in the TI film are $\pm e^{2}/2h$ and 0, respectively, leading to a half quantum mirror Hall effect by 1/21/21/2-1/2, serving as a metallic partner to the insulating quantum spin Hall effect. A brief proof for the half-quantization of a metallic band structure with considered symmetry constraints over the Fermi surface is also presented. Additionally, a field theoretical deduction for the half quantization, and a discussion on handling the Hall conductivity of a gapless Dirac fermion are provided.
@@
In the weak Zeeman field regime, the topological phases are characterized by focusing on $n=1$ matrix elements affecting the two gapless Dirac fermions near the surface. This framework clarifies the underlying physics behind the Chern insulator, axion insulator, and half QAHE, with symmetric, antisymmetric, or unilateral distribution of Zeeman fields at the surface of the TI film, respectively. The resulting Hall conductance exhibits quantized nature: $1+0$, $0+0$, and $1/2+0$ in units of $e^{2}/h$. Additionally, the mirror layer Chern number in the Chern insulator with symmetrically distributed magnetism is examined, revealing (1/4)(1/4)(1/2)(1/2)$(1/4)$ partition for the non-trivial band and (c/4)(c/4)(c/2)(-c/2)$(c/4)$ with $c\approx 1$ for the trivial band.
@@ II Massless and massive Dirac fermions in a topological insulator film
Figure 1: Schematic momentum dependent mass configurations (upper panel) and corresponding band structure of Dirac fermions (lower panel). The quantization of Hall conductivity is denoted by NN for half quantization in a metallic band and $C$ for quantization at the bottom of a gapped band. The colors assigned to the Dirac cones represent the sign of Berry curvature with red for positive and blue for negative. (a) On the left panel, two gapless Dirac fermions are shown, whose masses are zero at low-energy near the Dirac point (assumed to be $k=0$), while non-vanishing at high-energy with opposite signs, which we define as the chirality $\chi$ of a 2D gapless Dirac fermion. Such chirality unambiguously determines the sign of the loop integral of Berry connection around the Fermi surface, consequently determining the sign of half-quantized Hall conductivity. (b) In the middle, two trivially gapped massive Dirac fermions are present, with masses being either positive or negative for all $k$, leading to a sign change of Berry curvature and a totally vanishing Hall conductivity labeled by zero Chern number. (c) On the right panel, two non-trivial gapped Dirac cones are displayed, and the corresponding masses exhibit kink configurations with sign change between low and high-energy areas. Such non-trivial mass configuration indicates overall Berry curvature sign convergence, and leads to a non-vanishing Hall conductivity labelled by an integer Chern number. The non-triviality is also addressed by formally drawing states connecting conduction and valence bands, well-known as chiral edge states for a Chern band under open boundary conditions[13, 14].
@@
for the continuum and lattice model, respectively. Here we adopt a homogeneous in-film-plane parameter set with $a$ and $\lambda_{\parallel}$ as the in plane lattice constant and Fermi velocity, and $\bm{k}=(k_{x},k_{y})$ is the in-film-plane wavevector. Notice that an infinitely direct summed Dirac fermions exist in the continuum model, while there are only $2L_{z}$ species with $L_{z}$ the layer number along opened $z$-direction of the film in the lattice model. For the mainly concerned individual Dirac cone with a single Dirac point at $k=0$, its topological property is revealed based on a general discussion over the nature of its Hall conductivity quantization, as revealed in the schematic diagram Fig. 1. Especially, in the strong TI film with a single Dirac cone at $\Gamma$, aka $k=0$ point, the gapless pair of Dirac fermions carry $\pm e^{2}/2h$, as half-quantized Hall conductivity, while the gapped series are all trivial.
@@ II.1 The continuum model
$\varphi^{n}(s)$ $=C\begin{pmatrix}-is\lambda_{\perp}f_{+}^{n}\\ t_{\perp}\eta^{n}f_{-}^{n}\end{pmatrix},\penalty 10000\ E=m_{n},$ (4a)
$\chi^{n}(s)$ $=C\begin{pmatrix}t_{\perp}\eta^{n}f_{-}^{n}\\ is\lambda_{\perp}f_{+}^{n}\end{pmatrix},\penalty 10000\ E=-m_{n},$ (4b)
where the dependence on $(\bm{k},z)$ is inherited inside even/odd parity functions $f_{\pm}^{n}(\bm{k},z)$ and real factor $\eta^{n}(\bm{k})$, whose definition can be found in Appendix A. The $k$-dependent eigenvalue of $h(s)$ is represented by $\pm m_{n}(\bm{k}),\penalty 10000\ n=1,2,\cdots$, as a mass term, which can be solved in a closed manner through equations
@@
$\xi_{\alpha}$ $=\sqrt{-\frac{F}{D}+(-1)^{\alpha-1}\frac{\sqrt{R}}{D}},\penalty 10000\ \alpha=1,2,$ (5b)
@@
Figure 2: The momentum-dependent mass of Dirac fermions in a TI film as a continuum model. The lowest several momentum-dependent $m_{n}(\bm{k})$ along $k_{x}$ solved from closed equations Eq. (5) are presented, while the homogeneous in-plane nature of the model ensures that the asymptotic behavior of $m_{n}(\bm{k})$ is the same as $k\rightarrow\infty$. Here the film thickness $L=(L_{z}+1)c$ with $L_{z}=10$ is chosen here as a TI film with 1010 layers. The index $n$ is assigned such that $|m_{n}|$ increases with $n$. Especially notice the sign-jump behavior that $\mathrm{sgn}(m_{n}(\infty))=(-1)^{n}$. From here on, the model parameters on lattice for numerical calculations and verifications are set as $\lambda_{\parallel}=0.41\penalty 10000\ $\mathrm{e}\mathrm{V}$\$ eV, $\lambda_{\perp}=0.44\penalty 10000\ $\mathrm{e}\mathrm{V}$\$, $t_{\parallel}=0.566\penalty 10000\ $\mathrm{e}\mathrm{V}$\$, $t_{\perp}=0.4\penalty 10000\ $\mathrm{e}\mathrm{V}$\$, $m_{0}=0.28\penalty 10000\ $\mathrm{e}\mathrm{V}$\$, $a=b=1\penalty 10000\ $\mathrm{n}\mathrm{m}$\$, $c=0.5\penalty 10000\ $\mathrm{n}\mathrm{m}$\$ if with no specific indication[54]. Generically, these parameters can be determined through the first-principle calculations, and the specific choice here is for the sake of illustration. This parameter choice makes the bulk 3D TI a strong one with a single Dirac point at $\Gamma$. And for the continuum model discussed here, the substitution $\lambda_{\parallel}\rightarrow\lambda_{\parallel}a$ $\lambda_{\perp}\rightarrow\lambda_{\perp}c$, $_{\parallel}\rightarrow t_{\parallel}a^{2}, tt$$_{\perp}\rightarrow t_{\perp}c^{2}$ should be recognized.
@@
with $\gamma=m_{0}t_{\perp}/\lambda_{\perp}^{2},u=\lambda_{\perp}/2t_{\perp}$. The numerical result is shown in Fig. 3, with excellent agreement between the lowest order approximated gap and that from solving the set of non-linear equations, especially for relatively large $L$. The exponentially decaying tendency is best revealed by the logarithmic absolute value of mass gap at $k=0$, as its center decreases linearly with thickness, while the oscillating nature is revealed by the dips, which will extend to negative infinity at strict gap closing point, and the mass gap will reverse its sign before and after the dip, as shown directly by the $m_{1}(0)-L$ diagram and the inner amplified picture. Since $m_{1}(\infty)=m_{0}(\infty)<0$ is certain, we see that the oscillating behavior of $m_{1}(0)$ with thickness $L$ can drive $m_{1}(\bm{k})$ to share configuration that jumps between the one shown in Fig. 1(b) and (c), i.e., between a trivial band and a band with unit Chern number. Then for an ultra-thin film which owns two copies $\pm m_{1}(\bm{k})$ reflected by $\tau_{z}$ in Eq. (7), the $\mathbb{Z}_{2}$ topological index shows jumping behavior between $\mathbb{Z}_{2}=0$ and $\mathbb{Z}_{2}=1$, i.e., between a band insulator and a quantum spin Hall insulator[40, 59, 60, 41, 42]. We will not discuss further about this phenomenon except for giving an explicit $\mathbb{Z}_{2}(L_{z})$ oscillating diagram below in the lattice model subsection shown in Fig. 5. We emphasize here that the exponentially decaying gap will not affect physically observable topological phase, either for an insulating or metallic one, for a TI film with enough thickness.
@@ II.2 The lattice model
where energy and hopping matrices are $\mathcal{M}_{0}=(m_{0}-2\sum_{\mu}t_{\mu})\beta$, $\mathcal{T}_{\mu}=t_{\mu}\beta-i\frac{\lambda_{\mu}}{2}\alpha_{\mu}$, with $l$ and $\mu$ denoting site locations and three spatial directions, while $\{\beta,\alpha_{\mu}\}$ denoting Dirac matrices under standard Dirac representation $\beta=\sigma_{0}\tau_{z},\penalty 10000\ \alpha_{\mu}=\sigma_{\mu}\tau_{x},$ where Pauli matrices $\sigma_{\mu}$ and $\tau_{\mu}$ represent different degrees of freedom, respectively. For instance, one could choose them to represent spin and pseudo-spin (like orbital) ones. $\Psi_{l}$ represents vectorized Fermionic operator at site $l$. Notice that when adopting a full Fourier transformation upon all three spatial dimensions, i.e., an infinite bulk system, the Hamiltonian is transformed into the standard modified Dirac’s equation[27] on lattice $\mathcal{H}_{TI}=\sum_{\bm{k}}\Psi_{\bm{k}}^{\dagger}H(\bm{k})\Psi_{\bm{k}}$ where
@@
$\sin\xi_{\alpha}=\sqrt{1-\cos\xi_{\alpha}^{2}},\penalty 10000\ \alpha=1,2.$ (22)
Now, different from the continuum model, the set of equations give $L_{z}$ solutions $m_{n}(\bm{k}),n=1,2,\cdots,L_{z}$ including one surface state and $L_{z}-1$ purely trivial bulk states, if within suitable choice of parameters. This is essentially because now the Dirac equation is put on lattice, and the number of solutions is constrained by finite lattice constants. And the other set of $L_{z}$ masses are just the chiral partners with eigenvalues $m_{n}(\bm{k})$.
Figure 4: The momentum-dependent mass of Dirac fermions in a TI film on the lattice. Namely, $m_{n}(\bm{k}),\penalty 10000\ n=1,2,\cdots,L_{z}$ along $k_{x}$ are solved from closed equations Eq. (20) of lattice model with $L_{z}=40$. Again, index $n$ is assigned in the way that $|m_{n}(\pi,\pi)|$ increases with $n$. Especially notice the sign-jump behavior that $\mathrm{sgn}(m_{n}(\pi,\pi))=(-1)^{n}$.
@@
with $\chi=\pm$ labelling the mirror eigenvalue[33]. An example of $m_{n}(k)$ with $L_{z}=40$ is presented in Fig. 4. Among these Dirac fermions, two of them with $\pm m_{1}(\bm{k})$ are gapless Dirac cones with their low-energy states localized at top and bottom surfaces, while emerging into the bulk at their high-energy away from Dirac point, and the remaining fermions are all gapped. Notice that the same arguments about the projection as a trivial local unitary transformation and Heaviside Theta function form of the lowest solution (see below) can be made here, as in the continuum model.
@@ III.1.1 Gapless/metallic case
$h_{1,\chi}=\lambda_{\parallel}(\bm{k}\cdot\bm{\sigma})+\chi\Theta(-m_{0}(k))m_{0}(k)\sigma_{z},\penalty 10000\ \chi=\pm.$ (32)
@@ III.1.2 Insulating case
for a Dirac cone, where $d_{\min}=\min(d(k))$ denotes the bound of the global gap. Clearly, $\sigma_{H}/(e^{2}/h)=0,\pm 1$ appears, notifying trivial or non-trivial phases depending on the relative signs of low and high-energy masses, with the $\pm 1$ cases identified as the Chern insulator or equivalently, the quantum anomalous Hall effect. In our equivalent model, one sees from Fig. 2 that all $n\geq 2$ masses contains the same sign, and the corresponding Dirac cones are all trivial. And we come back to the statement that in a TI film, there are two gapless Dirac fermions with opposite half-quantized Hall conductance, while all other bands form paired trivial massive Dirac fermions. The quantized nature of the Hall conductance in insulating system, $\sigma_{H}=-Ce^{2}/h$, is referred to by the famous TKNN theorem[10], with its robustness against continuous non-gap-closing perturbations rooted in the topological nature of $C$ as the Chern invariant[69, 70].